Tutored Practice Problem 15.4.5 colos.com Use the graphical method to determine reaction order and rate constant. Concentrati please help
Answer #1

For first order reaction;

ln [At] = - k*t + ln [A0]

y = mx + c

Slope = m = -k = -2.43*10^-2 s^-1

Intercept (c) = ln[A0] = -3.86

A0 = e^(-3.86) = 0.021068

Thus, reaction is first order in PH3 .....Answer

Rate constant (k) = 2.43*10^-2 s^-1 = 0.0243 s^-1 .....Answer

Let me know if any doubts.